# Discovering novel order parameters in the Potts model: A bridge between machine learning and critical phenomena

### v1

@@ IV Conclusion
Through systematic finite-size scaling analyses, we have discovered and validated a family of alternative order parameters for the square-lattice $q=3$ and $q=4$ Potts models. These include: (1) the conventional order parameter based on dominant spin state occupancy, (2) novel parameters derived from secondary and minimal spin states, and (3) magnetization-like quantities emerging from reduced representations with only two spin components. We have rigorously established the intrinsic relationships between these alternative order parameters, demonstrating that they collectively provide multiple consistent characterizations of critical behavior. This unified framework explains the remarkable generalizability of Ising-trained machine learning models to Potts systems. The models remain effective because criticality is redundantly encoded in these various order parameters, persisting even when explicit spin state information is partially lost through representation reduction.
Our findings suggest these novel order parameters generalize to other $q$-state Potts models, different lattice geometries (e.g., triangular or 3D), and systems with extended interactions (anti-ferromagnetic couplings, next-nearest neighbors). They may further enable studies of real-time dynamics or non-equilibrium phase transitions. Crucially, this work bridges machine learning’s empirical success in phase classification with fundamental thermodynamic quantities, demonstrating that conventional order parameters are not unique descriptors of critical phenomena. Future work could explore these generalizations systematically or analyze how neural networks leverage such order parameters across diverse spin systems.
@@ Acknowledgments
Y.D. thanks Ke Xu for the assistance with the usage of HPC in the cluster. This work is supported by the University of Bergen through “Akademiaavtalen” (K.T. and N.S.), the Taishan Scholars Program under Grant No. TSQNZ20221162, and the Shandong Excellent Young Scientists Fund Program (Overseas) under Grant No. 2023HWYQ-106 (Y.D.). Y. D. thanks the support from the cluster in FIAS and the Norwegian e-infrastructure UNINETT Sigma2 with Project Nos. NS9753K and NN9753K for the HPC resources and data storage in Norway. The work of N.S. was done at FIAS, supported by the AI grant at FIAS of SAMSON AG.
Appendix A Finite-size scaling of alternative order parameters
A.1 $q=3$ Potts model on a square lattice
Figure A1: (Color online) (Left) Order parameter $M_{\mathrm{map}}$ as a function of temperature $T$ for the $q=3$ Potts model on a square lattice with six different system sizes $L$. (Middle) Scaled order parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ versus $T$, with curves intersecting near the theoretical critical temperature $T_{c}$, marked by the red vertical dashed line. (Right) Scaled order parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ against the rescaled temperature $tL^{1/\nu}$, demonstrating data collapse consistent with finite-size scaling.
@@ A.2 $q=4$ Potts model on a square lattice
Figure A8: (Color online) (Left) Order parameter $M_{\mathrm{map}}$ as a function of temperature $T$ for the $q=4$ Potts model on a square lattice with six different system sizes $L$. (Middle) Scaled order parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ versus $T$, with curves intersecting near the theoretical critical temperature $T_{c}$, marked by the red vertical dashed line. (Right) Scaled order parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ against the rescaled temperature $tL^{1/\nu}$, demonstrating data collapse consistent with finite-size scaling.

### v2

@@ IV Conclusion
Through systematic finite-size scaling analyses, we have discovered and validated a family of alternative (quasi)order parameters for the square-lattice $q=3$ and $q=4$ Potts models, grounded in the underlying principles of universality and symmetry. These include (1) the conventional order parameter based on dominant spin state occupancy, (2) order parameters derived from secondary and minimal spin states, and (3) magnetization-like quasiorder parameter emerging from reduced representations with only two spin components. We have rigorously established the intrinsic relationships between these alternative (quasi)order parameters, demonstrating that they collectively provide multiple consistent characterizations of critical behavior associated with different level of standard uncertainty. This unified framework explains the remarkable generalizability of Ising-trained machine-learning models to Potts systems. The models remain effective because criticality is redundantly encoded in these various (quasi)order parameters, persisting even when explicit spin state information is partially lost through representation reduction.
Motivated by these findings, we further note that other quantities, such as $|N_{1}/N-1/q|$, may also serve as (quasi)order parameters. This corresponds to an even more drastic mapping in which all except one of the Potts states are mapped to zero. We find that any quantity, such as $M_{\mathrm{map}}$ and $M_{1}\propto|N_{1}/N-1/q|$, that treats ordering in a subset of the Potts states differently from ordering in the other Potts states will have non-thermal fluctuations that increase as $T$ decreases and may obscure a peak in the thermal fluctuations near the critical point.
Our findings suggest that these (quasi)order parameters generalize to other $q$-state Potts models, different lattice geometries (e.g., triangular or 3D), and systems with extended interactions (anti-ferromagnetic couplings, next-nearest neighbors). They may further enable studies of real-time dynamics or non-equilibrium phase transitions. More broadly, this work provides a concrete example of how reduced spin representations can preserve the essential thermodynamic information required for identifying criticality, highlighting that conventional order parameters are not unique descriptors of critical phenomena and pointing toward the possibility that analogous quasiorder parameters, and the representations implicitly used by machine-learning models, may arise in a wider class of interacting spin systems.
@@ Acknowledgments
Y.D. thanks Ke Xu for the assistance with the usage of HPC in the cluster. This work was supported by the University of Bergen through “Akademiaavtalen” (K.T. and N.S.), the Taishan Scholars Program under Grant No. TSQNZ20221162, and the Shandong Excellent Young Scientists Fund Program (Overseas) under Grant No. 2023HWYQ-106 (Y.D.). Y.D. thanks the support from the cluster in FIAS and the Norwegian e-infrastructure UNINETT Sigma2 with Projects No. NS9753K and No. NN9753K for the HPC resources and data storage in Norway. The work of N.S. was done at FIAS, supported by the AI grant at FIAS of SAMSON AG.
Appendix A Finite-size scaling of alternative (quasi)order parameters for $q=3$ and $q=4$ Potts model on a square lattice
In this Appendix, we present additional finite-size scaling analyses for the alternative (quasi)order parameters introduced in the main text for the $q=3$ and $q=4$ Potts models on a square lattice. These results complement the main discussion by providing detailed numerical evidence for the scaling behavior with known theoretical or numerical critical temperatures and critical exponents, thereby supporting the robustness of our conclusions.
Figure A1: (Color online) (Left) Order parameter $M_{\mathrm{max}}$ as a function of temperature $T$ for the $q=3$ Potts model on a square lattice with six different system sizes $L$. (Middle) Scaled order parameter $M_{\mathrm{max}}L^{\Delta_{\sigma}}$ versus $T$, with curves intersecting near the theoretical critical temperature $T_{c}$, marked by the red vertical dashed line. (Right) Scaled order parameter $M_{\mathrm{max}}L^{\Delta_{\sigma}}$ against the rescaled temperature $tL^{1/\nu}$, demonstrating data collapse consistent with finite-size scaling.
Figure A2: (Color online) (Left) Quasiorder parameter $M_{\mathrm{map}}$ as a function of temperature $T$ for the $q=3$ Potts model on a square lattice with six different system sizes $L$. (Middle) Scaled quasiorder parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ versus $T$, with curves intersecting near the theoretical critical temperature $T_{c}$, marked by the red vertical dashed line. (Right) Scaled quasiorder parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ against the rescaled temperature $tL^{1/\nu}$, demonstrating data collapse consistent with finite-size scaling.
@@
Figure A9: (Color online) (Left) Quasiorder parameter $M_{\mathrm{map}}$ as a function of temperature $T$ for the $q=4$ Potts model on a square lattice with six different system sizes $L$. (Middle) Scaled quasiorder parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ versus $T$, with curves intersecting near the theoretical critical temperature $T_{c}$, marked by the red vertical dashed line. (Right) Scaled quasiorder parameter $M_{\mathrm{map}}L^{\Delta_{\sigma}}$ against the rescaled temperature $tL^{1/\nu}$, demonstrating data collapse consistent with finite-size scaling.